Current Attempt in Progress In 2022, Bramble Corp. had cash receipts of $54880 and cash disbursements of $31360. Brambles caView Policies Current Attempt in Progress In 2022, Sheridan Company had cash receipts of $29400, including additional owner iView Policies Current Attempt in Progress Blue Spruce Corp.s trial balance reported the following balances at the end of itsCurrent Attempt in Progress Suppose that Pharoah Company produced and sold 4800 laptop computers during 2022. It reported $16Current Attempt in Progress During January 2017, its first month of operation, Pharoah Company earned net income of $8430 and
Answer #1

1. Opening Balance= Closing balance+ Cash disbursement - Cash Receipt

76440+31360-54880= 52920

Answer: Option A :52920

2. Balance at January 2022 = Closing balance (December 2022)+ Cash disbursement+ Purchase of equipment - Cash Receipts - Additional owner Investment.

54600+16800+12000 - 29400 -10000= 44000

Answer = 44000

Note : Answer will be 44000. But there is no option in the above list.

3. Total credit balance will be the total liabilities items and revenue items . so

total credit= Accounts payable+ notes payable+ Revenue

= 3250+4870+33640= 41760

4. Free cash flow is the amount left over after meeting the expenditures.

161200 - 10660 - 17400= 148800

5. balance = Net income - dividend

8430- 2480= 5950 credit